Object A and Object B are charged with +47.4 μ coulombs and +16.3 μ coulombs, respectively. Given that the two objects are separated by 0.64 m, Which of the following statements is/are valid?





a. Object A attracts object B with a force of 27.1 N



b. Object A repels object B with a force of 17:0N



c. Object B attracts object A with a force of 27.1N



d. Object A attracts object B with a force of 118.7 N



e. Object 8 repeli object A with a force of 120 N



1. Object Battracts object A with a force of 118.7 N



Object 8 repels object A with a force of 433 N



n Object A repels object 8 with a force of 43. N

1
Expert's answer
2022-03-10T10:17:50-0500

We know that Electric

force

"F=\frac{kq_1q_2}{r^2}"


"F=\frac{9\times10^9\times47.4\times10^{-4}\times16.3\times10^{-13}}{0.64^2}=17N"

Repulsion force

Option (b) is correct option